OpenStudy (anonymous):

Let k and n be positive integers, with k < n. Write n! -- k! without using factorials; simplify your answer.

OpenStudy (anonymous):

\[\frac{n!}{k!}\]

hartnn (hartnn):

n! = n(n-1)(n-2)....(n+k-n+1)k! or you can take n-k= p as n>k n!= (p+k)! = (p+k)(p+k-1).....(p+k-(p-1))k! then put p= n-k